OpenStudy (anonymous):

A population of 100 contains 12 nonconforming items. a) What is the probability of selecting a random sample of 25 items containing 3 to 5 nonconforming items? Sampling is without replacement. b) Calculate the parameters mean and variance of the distribution of nonconforming items in samples of 25. c) Explain on the meaning of the quantities mean and variance in the context of this problem.

OpenStudy (anonymous):

did you get this from a test in online school

OpenStudy (anonymous):

no actually my teacher gave this question to practice before test.

OpenStudy (anonymous):

oh ok :) do you know where to start?

OpenStudy (anonymous):

i can't understand the way he explains that's why i posted this question so that i can learn how to do these types of problems

OpenStudy (anonymous):

well duh, unfortunately all i can help with is a way to guide you through it because i do not understand this either :/

OpenStudy (anonymous):

@BlackLabel you say something or stop veiwing

OpenStudy (anonymous):

hahahaah sorry just negotiating a business deal but I can be of assistance

OpenStudy (anonymous):

thanks guys for helping me. i really appreciate it

OpenStudy (anonymous):

no problem i will help the best i can :)

OpenStudy (anonymous):

so do you know how tis problem is supposed to be worked out ?

OpenStudy (anonymous):

Ok firstly we need to use the binomial distribution. You familiar with that?

OpenStudy (anonymous):

i think is it that the one with p(xabsolute N,D, n) formula

OpenStudy (anonymous):

it think this equation is called hypergeometric distribution

OpenStudy (anonymous):

Dinner time ill bbl. Try to get someone else to help u Ohhh Ya that may be a better option. Ill check back in a half hr

OpenStudy (anonymous):

ok

OpenStudy (kropot72):

\[P(3\ nonconforming)=\frac{\left(\begin{matrix}12 \\ 3\end{matrix}\right)\left(\begin{matrix}88 \\ 22\end{matrix}\right)}{\left(\begin{matrix}100 \\ 25\end{matrix}\right)}\]

OpenStudy (anonymous):

so x is 3 and n is 25 and defectives are 12 right. i did not get 1 thing that in the question it says 3 to 5 nonconforming so we can assume one?

OpenStudy (anonymous):

i think i got how to do part a i got the answer 0.2750

OpenStudy (kropot72):

The hypergeometric distribution is generated when a sample of size n is taken, without replacement) from a population containing a items of one particular type and b of others. The probability that the sample contains exactly x of the particular type is given by the relevant term in the hypergeometric distribution. \[p(x)=\frac{\left(\begin{matrix}a \\ x\end{matrix}\right)\left(\begin{matrix}b \\ n-x\end{matrix}\right)}{\left(\begin{matrix}a+b \\ n\end{matrix}\right)}\]

OpenStudy (kropot72):

@Gurmeet Good work! Your answer to part a is correct.

OpenStudy (anonymous):

great :)

OpenStudy (anonymous):

i didn't get it how to do part b and c

OpenStudy (anonymous):

could you also help me do these as well?

OpenStudy (anonymous):

you explain really good

OpenStudy (kropot72):

b. For the hypergeometric distribution: \[mean=\frac{na}{a+b}\] \[variance=\frac{nab(a+b-n)}{(a+b)^{2}(a+b-1)}\]

OpenStudy (anonymous):

is mean 3

OpenStudy (kropot72):

Yes, your result for the mean is correct.

OpenStudy (anonymous):

and is the variance 2

OpenStudy (kropot72):

Correct! The variance is 2.

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

what does part c means i didn't get it

OpenStudy (kropot72):

The sample mean is the most probable number of of nonconforming items in a sample. The sample variance 2 is a measure of variation and is the square of the standard deviation. In this case: \[Sample\ variance,\ s ^{2}=2\]

OpenStudy (kropot72):

*number of nonconforming items

OpenStudy (anonymous):

ok so basically we are just explaining our solution right ?

OpenStudy (kropot72):

Yes, we are explaining what the sample mean and the sample variance actually mean regarding the question.

OpenStudy (anonymous):

Out of 10 loaves on a shelf, 4 are three days old. If the store manager selects 5 loaves at random, what is the probability that: a) one of the loaves is three days old? Answer: 0.2381 b) x of the loaves is three days old? Answer: \[P(x of the loaves) = (4 nCr x) (6 nCr (5-x)) / (10nCr5)\] Did i do this correctly? i used hypergeometric distrubution to solve it

OpenStudy (kropot72):

Yes, your answers to a) and b) are correct.